OpenStudy (anonymous):

I am 14 years old and I couldn't solve this trigonometry question, if tan2x=3/4 then what is the value of sinx?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

Tan 2x = 3/4 = .75 2x = arc tan(.75) = 36.87 degrees x= 36.87/2 = 18.44 degree sinx= sin 18.44 = 0.31622 use calculator

OpenStudy (anonymous):

did u get the process of doing it korcan?

OpenStudy (anonymous):

well thanky you , but I found the answer as sinx= 1/3√5, and whats the meaning of arc?

OpenStudy (anonymous):

I don2T know if my answer is true.

OpenStudy (anonymous):

14yrs old?

OpenStudy (anonymous):

that makes you like yr9 , this is a yr11/12 question

OpenStudy (anonymous):

mark o. you did it wrong , they want an exact value for it

OpenStudy (anonymous):

you need to use double angle expansion

OpenStudy (anonymous):

YEa i am 14 years old and i learned trgionometry and function when i was 12 ^^

OpenStudy (anonymous):

\[\tan(2x) = \frac{2\tan(x)}{1-\tan^2(x)} \]

OpenStudy (anonymous):

mhmh just have a look at my answer mybe?

OpenStudy (anonymous):

and check it please?

OpenStudy (anonymous):

well u need to put the answer right awy so that we know tht u want to arrive at the answer w/c is 1/3sqrt5.....lol

OpenStudy (anonymous):

1/(3sqrt5) = 0.316227766 using a calculator... does your teacher allow you to use calculator? ...lol...

OpenStudy (anonymous):

you need to use a scientific calculator in order to solve trig, sometimes calculator serves you to check if your answer is correct....

OpenStudy (anonymous):

im sure you can use a table of trig fuctions?..lol

OpenStudy (anonymous):

arctangent is like this: tan(36.87deg)=3/4 now the reverse of it is.... 36.87deg=arc tan(3/4) or arc tan(3/4) = 36.87deg

OpenStudy (anonymous):

did you get it korcan?

OpenStudy (anonymous):

you can ask question if you are not sure so that we can explain to you better..lol

OpenStudy (anonymous):

OKA i got it thank you ^^but my teacher doesn2T let me use calculater, :( or table of trig fuctions :( but i know how to use it ^^

OpenStudy (anonymous):

and i also dun have a calculator at hause xd ^^ and dun have calculator on p ctoo but i can program a calculator when i need ^^